ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 156, 2006

Alesker, S. Theory of valuations on manifolds, I. Linear spaces  311

Artstein-Avidan, S. A Bernstein–Chernoff deviation inequality, and geometric properties of random families of operators

187

Artstein-Avidan, S., O. Friedland, V. Milman and S. Sodin Polynomial bounds for Large Bernoulli sections of ℓ_{1}^{N}

141

Benanti, F. and I. Sviridova Asymptotics for Amitsur’s Capellitype polynomials and verbally prime PI-algebras

73

Bobkov, S. G., C. Houdré and P. Tetali The subgaussian constant and concenentration inequalities

255

Cheng, Lixin Ball-covering property of Banach spaces  111

Danilenko, Alexandre I. Mixing rank-one actions for infinite sums of finite groups

341

Dettweiler M. and S. Wewers Variation of local systems and parabolic cohomology

157

Dow, A. Two to one images and PFA  221

Downarowicz, T. Minimal models for noninvertible and not uniquely ergodic systems

93

Ferenczi, V. Minimal subsurfaces and isomorphically homogeneous sequences in Banach space

125

Friedland, O. See S. Artstein-Avidan

Golberg, A. On generalization of Menshoff’s theorem  243
Han, Pigong. Multiple solutions to singular critical elliptic equations 359

Houdré, C. See S. G. Bobkov

Ito, N. and A. Mann Counting classes and characters of groups of prime exponent

205

Lu, Guangcun Gromov–Witten invariants and pseudo symplectic capacities

1

Mann, A. See N. Ito

Milman, V. See S. Artstein-Avidan

Perdomo, Oscar Metrics on products of surfaces with non-positive sectional curvature

65

Simon, R. S. Value and perfection in stochastic games  285

Sodin, S. See S. Artstein-Avidan Sviridova, I. See F. Benanti Tetali, P. See S. G. Bobkov Wewers, S. See M. Dettweiler